OpenStudy (anonymous):

What is the angle between the two hands of clock when it is 7 : 30?

OpenStudy (anonymous):

Should be 45 degrees, yeah? The minute hand is straight down, and the hour hand is between the 7 and the 8, which is 45 degrees away.

OpenStudy (anonymous):

(360)/(12*2)

OpenStudy (anonymous):

15deg

Parth (parthkohli):

When it is 7:30, the hour hand is not exactly at 7. It is in the middle of 7 and 8... 24 + 12 = 36

OpenStudy (anonymous):

Depends on whether the hour hand moves continuously or discretely. I answered for continuously.

OpenStudy (anonymous):

ya i made a mistake ! what i said is 6:30

OpenStudy (anonymous):

How did you get 45 @nbouscal

OpenStudy (anonymous):

Visually. 

OpenStudy (anonymous):

Those angles are not exact, obviously, but do you see what I mean? It's halfway to 90 degrees.

Parth (parthkohli):

24 means 1/12 * 360, and between it would be 1/24 * 360... 24 + 12 = 36

OpenStudy (anonymous):

can we calculate using maths logic

OpenStudy (anonymous):

@ParthKohli ans. is 45.... but I don't get it ..

OpenStudy (anonymous):

If we treat it as a unit circle on the xy-plane, with zero degrees at the 3 on the clock, the minute hand is at 270 degrees, and the hour hand is at 225 degrees.

OpenStudy (anonymous):

We can calculate that by knowing that each number on the clock is 30 degrees apart from each other (360/12). And we know that the hour hand is halfway between the 7 and the 8.

Parth (parthkohli):

Actually lol, 360/12 = 30, and 360/24 = 15 30 + 15 = 45

OpenStudy (anonymous):

Thanx @nbouscal

OpenStudy (anonymous):

great man.... ty @ParthKohli